Suppose a machine is set to fill a candy package with 56 candies per bag. A sample of 9 bags results in a mean of 57 and a sample standard deviation of 1.0. How many degrees of freedom are used to test the hypothesis?
Answer #1

Solution:-

Given that,

mean = = 57

standard deviation = = 1.0

n = 9

degrees of freedom = n - 1 = 9 - 1 = 8